OpenStudy (anonymous):

Sam and Harry are family. Sam is currently is three times Harry's age. Sam's age is also 10 more than twice Henry's age. What are their current ages? choices: Sam is 24 years old, and Harry is 8 years old. Sam is 39 years old, and Harry is 13 years old. Sam is 27 years old, and Harry is 9 years old. Sam is 30 years old, and Harry is 10 years old.

OpenStudy (anonymous):

@sleepyjess could you help me out?

OpenStudy (sleepyjess):

S = 3H S = 2H + 10

OpenStudy (sleepyjess):

set them equal to each other, 3H = 2H + 10, solve for H

OpenStudy (anonymous):

I dont understand on how to do that though :/

OpenStudy (anonymous):

@sleepyjess I dont understand how to do that though :/

OpenStudy (sleepyjess):

what don't you understand?

OpenStudy (anonymous):

how to solve for H

OpenStudy (sleepyjess):

Well, right now we are adding 2H, so subtract 2H from both sides.

OpenStudy (sleepyjess):

Subtracting 2H from both sides gives us 3H - 2H = 2H - 2H + 10

OpenStudy (anonymous):

So is that part what we get? and do we then add the 2H? Sorry I am having a hard time with this.

OpenStudy (sleepyjess):

what is 3H-3H?

OpenStudy (sleepyjess):

*3H-2H

OpenStudy (anonymous):

H?

OpenStudy (sleepyjess):

Yep, so H = 2H - 2H +10. What is 2H - 2H?

OpenStudy (anonymous):

H as well?

OpenStudy (anonymous):

correct @sleepyjess

OpenStudy (anonymous):

2H+10?

OpenStudy (sleepyjess):

2H - 2H is just 0

OpenStudy (anonymous):

so then it would be 0+10 or 0H+10?

OpenStudy (sleepyjess):

H = 0 + 10, H = 10

OpenStudy (anonymous):

The what do you do afterwards?

OpenStudy (sleepyjess):

Just plug in 10 for H into one of the equations and find S

OpenStudy (anonymous):

okay

OpenStudy (sleepyjess):

Or just choose the only answer where Harry is 10